Circular No. 6241

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ae IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of a supernova (V = 18.5) on two Technical Pan films obtained on Sept. 22.97 and 27.03 UT by D. Albanese and himself with the OCA Schmidt telescope; the supernova is located at α = 23^{h}16^{m}55^{s}.65, δ = −2^{o}04′36′′.4 (equinox 2000.0), which is 5′′.4 west and 0′′.5 north of the galaxy’s center (the galaxy itself being of mag B ∼ 17.0). Nothing is visible at this location on the Palomar Sky Survey prints or the Digital Sky Survey. A nearby star (V ∼ 15) has position end figures 17^{m}02^{s}.55, 05′28′′.9.

S. Benetti, European Southern Observatory (ESO), reports: “Inspection of a fully-reduced CCD spectrogram (range 374–690 nm, resolution 1.5 nm), obtained on Oct. 1.05 UT with the ESO 3.6-m telescope (+ EFOSC1) at La Silla, confirms this object as a type-Ia supernova ∼ 10 days after maximum. The spectrum is dominated by strong emission lines measured at 421.2, 448.1, 491.6, 527.0, and 560.3 nm. The S II 565.4-nm line is still present (absorption measured at 583.5 nm). From the Si II 635.5-nm absorption line (measured at 657.0 nm), an expansion velocity of ∼ 9850 km/s is derived (the redshift of the parent galaxy being 20 670 km/s, as measured from the interstellar Na D absorption blend, present in the supernova’s spectrum).”

GRO J1750–27 T. Dotani, R. Fujimoto, F. Nagase, and H. Inoue, Institute of Space

and Astronautical Science, on behalf of the ASCA team, communicates: “We observed the region of GRO J1750–27 (cf. IAUC 6222) with ASCA between Sept. 25.36 and 26.05 UT. The pulsar was detected near the edge of ASCA’s GIS field-of-view (and outside the SIS field-of-view), near the northwest corner of the BATSE error box (source position α = 17^{h}49^{m}12^{s}.0, δ = −26^{o}38′50′′; equinox 2000.0; uncertainty radius 2′). The average x-ray flux of the source was 6 × 10^{−10} erg cm^{−2} s^{−1} in the band 2–10 keV. We detected clear pulsation with a period of 4.452 ± 0.006 s (corrected for the solar-system barycenter), which confirms the source to be GRO J1750–27. The energy spectrum is approximated by a power law with a photon index of 0.8 and an absorption column of 2 × 10^{22} cm^{−2}.”

1995 October 2 Daniel W. E. Green